OpenStudy (anonymous):

I'm really not understanding how to solve this problem and this will be on my final. What is the solution of the system? 4x - 4y = -40 -4x + 7y = 55

hero (hero):

Hint: Add them together

hero (hero):

(4x - 4x) + (7y - 4y) = 55 - 40

hero (hero):

You can only add the equations together in cases where one of the variables will cancel. This is one of those cases.

OpenStudy (anonymous):

So I would get (5,5)

OpenStudy (anonymous):

y = x+10 I think.

OpenStudy (anonymous):

http://puu.sh/1zSbn

hero (hero):

@assaulth3ro911 no bro

hero (hero):

@AstralDnB They're not both positive. One coordinate is positive. The other, negative.

hero (hero):

You have to pay attention to signs.

OpenStudy (anonymous):

ok. I re-did everything and got (5,-5)

hero (hero):

Now you have it backwards

OpenStudy (anonymous):

Oh sorry about. Got excited and posted the answer backwards. Thnx for you help.